find the sum of three consecutive even integers such that four times first and three times the second integer is two less than six times the third integer. find the integers.
Answer by CubeyThePenguin(3113)

consecutive even integers: (x-2), x, (x+2)

4(x-2) + 3x = 6(x+2) - 2
4x-8 + 3x = 6x+12 - 2
7x - 8 = 6x + 10
x = 18

The numbers are 16, 18, and 20.